# The effects of dark energy on the matter-gravity coupling

### v1

@@ Abstract
Dark energy perturbations are expected to modify the evolution of cosmological perturbations, producing observable effects on scalar perturbations and gravitational waves. We show that the effects of dark energy for a generic modified gravity model can be encoded in a momentum, space and polarization dependent effective matter-gravity coupling, in agreement with the results obtained in some specific theories, and that this is not just related to the effective Planck mass appearing in the action as a coefficient of the Ricci’s scalar. Using appropriately defined effective energy tensors, we show that the method can be extended to include also the effects of higher order perturbations. The generality of this effective description of the effects of dark energy on cosmological perturbations gives a solid theoretical justification for the phenomenological approaches often used in the analysis of cosmological observations, allowing to understand the mapping between the results of phenomenological analysis of observational data and theoretical models. The momentum and polarization dependence of the effective gravitational coupling implies that the gravitational energy emitted by a binary coalescence can have an additional frequency and polarization dependence, due to the effects of dark energy perturbations.
@@ I Introduction
Dark energy accounts for most of the Universe energy density today, but its fundamental physical nature is still mysterious. For this reason modified gravity theories have been proposed as dark energy models, such as for example Horndeski Horndeski:1974wa. In these theories the field equations are modified with respect to general relativity (GR), and the new equations can be conveniently re-cast in the form of Einstein equations with a running Planck mass and effective stress-energy tensor associated to the new fields and geometrical terms, which is interpreted as dark energy.
This effective dark energy tensor has perturbations, which modify the GR cosmological perturbations equations, introducing new scale and time dependent effects. We show that these effects can be encoded in different effective time and momentum and polarization dependent matter-gravity couplings. As an application we discuss how this implies that the gravitational waves (GW) energy radiated by a binary system can have an additional frequency and polarization dependency, due to the interaction with dark energy.
@@ II Modified gravity theories
$\mathcal{L}_{\rm{tot}}=\sqrt{-g}\,\left(\frac{1}{2}M_{eff}^{2}R+\rm L_{DE}+\rm L_{m}\right)\,$, (1)
@@
$G_{\mu\nu}=\frac{1}{M_{\rm{eff}}^{2}}\left[T^{m}_{\mu\nu}+T^{DE}_{\mu\nu}\right]\,.$ (2)
@@ III Cosmological effects on scalar perturbations
When studying the effects of dark energy on the background, the energy momentum tensors ar normally treated separately, but for cosmological perturbations the effects of dark energy are often modeled in terms of a modification of the general relativity (GR) equations, in terms of an effective gravitational coupling Ishak:2024jhs.
For example, in the Newton gauge the linearized scalar perturbations equation with respect to a FRW background takes the form
@@
$G^{\Psi}_{\rm{eff}}=\frac{1}{M^{2}_{\rm eff}}\left(1+\frac{\delta\hat{\rho}_{DE}}{\delta\hat{\rho}_{m}}\right)\,.$ (5)
See for example DeFelice:2011hq for an explicit calculation of $G^{\Psi}_{\rm{eff}}$, based on analytical solutions of the perturbations equations for Honrdeski theories. Note that since the cosmological perturbations solutions depends on both space and time, also $G^{\Psi}_{\rm{eff}}$ is both space and time dependent. Given the generality of the assumptions made to derive eq.(26), we can conclude that an effective Poisson equation of the form given in eq.(26), can provide a general effective description of the effects of dark energy on structure formation.
@@
$\frac{k^{2}}{a^{2}}\Psi_{k}=\frac{1}{2}a^{2}G^{\Psi}_{\rm{eff}}(t,k)\delta\rho_{m,k}\\$ (6)
where we have now dropped the hat for ease of notation. This effective description allows to make the general prediction that the effective gravitational coupling can be both scale and time dependent, which can be useful for model independent analysis of observational data, treating $G^{\Psi}_{\rm{eff}}(t,k)$ as a phenomenological function constrained by data.
This effective approach requires to know the cosmological perturbations equations solutions to compute the effective gravitational coupling, i.e. it gives a general procedure to obtain the effective coupling, but it does not predict its explicit analytical form in terms of langrage coefficient. In this regard it is useful to use the effective field theory of dark energy to derive general analytical expression for $G^{\Psi}_{\rm{eff}}$, with the advantage of allowing to derive precise predictions for the interconnection with other observables Romano:2025apm; Romano:2025pcs.
A similar approach can also be applied to the other Bardeen potential $\Phi$, and for the equation relevant for weak lensing it gives
$\nabla^{2}(\Psi+\Phi)=a^{2}G^{\Psi+\Phi}_{\rm eff}\delta\rho_{m}\,\,$, (7)
which in momentum space takes the form
$\frac{k^{2}}{a^{2}}(\Psi_{k}+\Phi_{k})=a^{2}G^{\Psi+\Phi}_{\rm eff}(t,k)\delta\rho_{m,k}\,\,.$ (8)
Examples of explicit analytical calculations of $G^{\Psi+\Phi}$ and $G^{\Psi}$ for Horndeski theories in the deep sub-horizon limit, can be found in Linder:2015rcz. Note that this definition of effective gravitational coupling is consistent with the scale and time dependent ansatz used in the phenomenological analysis of large scale structure data Ishak:2024jhs.
@@ IV Cosmological effects on gravitational waves
where $h_{+}$ and $h_{\times}$ are the tensor modes, $h_{V1,2}$ the vector polarizations, and $h_{S}$ and $h_{L}$ the transverse and longitudinal scalar modes. Ignoring the interaction of GWs with matter, the free action is given by Romano:2024apw; Gubitosi:2012hu
@@
where $\mathcal{H}=a^{\prime}/a$. Including the interaction terms, we get
@@
Note that since cosmological perturbations are space and time dependent, in general also $G^{ij}_{\rm{eff}}(\eta,x)$ is expected to be space and time dependent. Contrary to GR, in a general modified gravity theory the vector and scalar degrees of freedom cannot be gauged away, because even in absence of matter perturbations the DE effective energy tensor will act as source for these polarizations. Note that the GWs propagation equation can be always re-cast in the form in eq.(12), by absorbing in $\delta T^{DE}_{ij}$ any modified gravity effect, including for example graviton mass terms. An alternative effective approach consists in encoding the effects of the source term in a frequency-time-polarization dependent effective speed Romano:2022jeh, which can be convenient for the calculation of the effects on gravity modification on the GW-EMW distance ratio.
@@ V General relativity limit
VI Encoding higher order perturbations in the effective coupling
@@
We have so far considered linear order metric perturbations, and at higher order $p$ other gauge invariant variables ${h}^{(p)}_{\mu\nu},\Psi^{p},\Phi^{p}$ can be defined Mollerach:2003nq; DeLuca:2019ufz; Chang:2020iji. Expanding the field equations (2) to higher orders in perturbations will give new equations which can always be re-cast in the canonical form Romano:2024apw
$\Box h^{(p)}_{ij}=\frac{a^{2}}{M^{2}_{\rm{eff}}}\left[\delta T^{m(p)}_{ij}+\delta T^{DE(p)}_{ij}\right]=a^{2}{G^{ij(p)}_{\rm{eff}}}\delta T^{m(p)}_{ij}\,$, (19)
[…]
This can always be achieved by appropriately defining the effective perturbed energy-stress tensor $\delta T_{ij}^{DE(p)}$, even if the d’Alambert operator does not appear explicitly in the perturbed equations, by adding it on both sides of the equations. The d’Alambert operator is defined w.r.t. to the same background metric in all these equations, so they can be summed to give
$\Box h^{(N)}_{ij}=a^{2}{G^{ij(N)}_{\rm{eff}}}\delta T^{m(N)}_{ij}\,$, (21)
where we have used the fact that the d’Alambert operator is linear, and defined the summed GWs and energy-tensor perturbations
$h^{(N)}_{ij}=\sum^{N}_{p=1}h^{(p)}_{ij}\quad,\quad T^{(N)m}_{ij}=\sum^{N}_{p=1}\delta T^{(p)m}_{ij}\,$ (22)
and
$G^{ij(N)}_{\rm{eff}}=\left[\sum^{N}_{p=1}\frac{\delta T^{(p)m}_{ij}}{G^{ij(p)}_{\rm{eff}}}\right]^{-1}\delta T^{(N)m}_{ij}$ (23)
For the modified Poisson’s equation we can obtain a similar result, by recasting the perturbed equations of order p in the canonical form
$\nabla^{2}\Psi^{(p)}=a^{2}G^{\Psi(p)}_{\rm{eff}}\delta{\rho}_{m}\,$,\\ (25)
@@
$G^{\Psi(N)}_{\rm{eff}}=\left[\sum^{N}_{p=1}\frac{\delta\rho^{(p)}_{m}}{G^{\Psi(p)}_{\rm{eff}}}\right]^{-1}\delta\rho^{(N)}_{m}\,$. (29)
@@
VII Physical implications for scalar and tensor perturbations
We have shown that the effects of dark energy on cosmological perturbations can be modeled by a momentum, time and polarization dependent effective gravitational coupling. A direct implication of this general result is that the gravitational energy radiated by a binary system can acquire an additional dependence on the GWs frequency and polarization. This is due to the fact that the calculation of the radiated power involves the solution of the wave equation for a binary coalescence, and the momentum and polarization dependence of the effective gravitational coupling will induce new effects. Formally we can write the solution of the wave equation 19 at first order in perturbations as
$h_{ij}=\Box^{-1}\left({G^{ij}_{\rm{eff}}}\delta T^{m}_{ij}\right)$ (30)
which differ from the GR case in which $G^{ij}_{\rm{eff}}$ is a constant, and where we have assumed a Minkowski background. Since $G^{ij}_{\rm{eff}}$ is polarization and space dependent, different polarizations can have a different radiated power. Writing the wave equation in momentum space, the momentum dependence of $G^{ij}_{\rm{eff}}$ would also induce a new additional frequency dependency. In a general modified gravity theory there can be GWs of scalar and vector polarization, which cannot be gauged away, contrary to GR. Performing a scalar-vector-tensor decomposition of the metric perturbations, for the tensor part, i.e. the transverse and traceless (TT) piece Flanagan:2005yc, at leading order we obtain
$h^{TT}_{ij}=\frac{2G^{ij}_{\rm{eff}}}{r}\ddot{Q}^{TT}_{ij},$ (31)
For scalar perturbations the scale dependence of the effective gravitational coupling allows to model phenomena such as the screening mechanism, i.e. a suppression the effects of dark energy on small scales, in order to pass local gravitational tests.
VIII Conclusions
For scalar perturbations the scale dependency of the effective gravitational coupling allows provide an effective description for the screening mechanism, i.e. the suppression of the effects of dark energy on small scales, which is often necessary in order to pass local gravitational tests.

### v2

@@ Abstract
Dark energy perturbations are expected to modify the evolution of cosmological perturbations, producing different observational effects. We show that these effects can be encoded in a momentum, space and polarization dependent effective matter-gravity coupling.
For scalar perturbations the effective gravitational coupling can be locally negative, in regions with local dark energy under-densities. For adiabatic perturbations the effective gravitational coupling for scalar perturbations can be negative for phantom dark energy, providing an explanation for the observed structure suppression at low redshift, consistent with other independent evidences of evolving dark energy.
@@ I Introduction
Dark energy accounts for most of the Universe energy density today, but its fundamental physical nature is still mysterious. For this reason modified gravity theories have been proposed as dark energy models, such as for example Horndeski gravity Horndeski:1974wa, or other extensions Gleyzes:2014qga. In these theories the field equations are modified with respect to general relativity (GR), and the new equations can be conveniently re-cast in the form of Einstein equations with a running Planck mass and an effective stress-energy tensor associated to the new fields and geometrical terms, which is interpreted as an effective dark energy stress-energy tensor. This effective dark energy tensor has perturbations, which modify the GR cosmological perturbations equations, introducing new scale and time dependent effects. We show that these effects can be encoded in different effective time, momentum and polarization dependent matter-gravity couplings. This implies for example that the gravitational waves (GW) energy radiated by a binary system can have an additional frequency and polarization dependency, due to the interaction with dark energy.
For scalar perturbations the effective gravitational coupling can be locally negative in regions with local dark energy underdensities. For scalar adiabatic perturbations the scalar effective gravitational coupling can be negative for background phantom dark energy DESI:2024mwx, providing an explanation for the observed structure suppression at low redshift Chen:2024vuf, and a consistency test between the background and perturbations effects of dynamical dark energy.
@@ II Modified gravity theories
$\mathcal{L}_{\rm{tot}}=\sqrt{-g}\,\left(\frac{1}{2}M_{\rm{eff}}^{2}R+\rm L_{DE}+\rm L_{m}\right)\,$, (1)
@@
$G_{\mu\nu}=\frac{1}{M_{\rm{eff}}^{2}}\left[T^{m}_{\mu\nu}+T^{DE}_{\mu\nu}\right]\,$, (2)
@@ III Cosmological effects on scalar perturbations
When studying the effects of dark energy on the background, the energy momentum tensors are normally treated separately, but for cosmological perturbations the effects of dark energy are often modeled in terms of a modification of the general relativity (GR) equations, in terms of an effective gravitational coupling Ishak:2024jhs.
For example, in the Newton gauge the linearized scalar perturbations equation with respect to a Friedmann-Robertson-Walker (FRW) background takes the form
@@
$G^{\Psi}_{\rm{eff}}(\eta,x^{i})=\frac{1}{M^{2}_{\rm eff}}\left(1+\frac{\delta\hat{\rho}_{DE}}{\delta\hat{\rho}_{m}}\right)\,.$ (5)
The effective eq.(4) admits by construction $\hat{\psi}$ as a solution, so it provides an effective equivalent description of the system, which can always be obtained form the solutions of the perturbations equations. See for example DeFelice:2011hq for an explicit calculation of $G^{\Psi}_{\rm{eff}}$, based on analytical solutions of the perturbations equations for Honrdeski theories. Note that this effective description is only valid for $\delta\rho_{m}\neq 0$. Since the cosmological perturbations solutions depends on both space and time, also $G^{\Psi}_{\rm{eff}}$ is both space and time dependent. Given the generality of the assumptions made to derive eq.(4), we can conclude that an effective Poisson equation of the form given in eq.(4), can provide a general effective description of the effects of dark energy on structure formation.
@@
$\frac{k^{2}}{a^{2}}\Psi_{k}=-\frac{1}{2}a^{2}G^{\Psi}_{\rm{eff}}(\eta,k)\delta\rho_{m,k}\\$ (6)
where we have defined
$G^{\Psi}_{\rm{eff}}(\eta,k)=\frac{1}{M^{2}_{\rm eff}}\left(1+\frac{\delta\rho_{DE,k}}{\delta\rho_{m,k}}\right)\,$, (7)
and we have now dropped the hat for ease of notation. This effective description allows to make the general prediction that the effective gravitational coupling can be both scale and time dependent, which can be useful for model independent analysis of observational data, treating $G^{\Psi}_{\rm{eff}}(t,k)$ as a phenomenological function constrained by data.
This effective approach requires to know the cosmological perturbations equations solutions to compute the effective gravitational coupling, i.e. it gives a general procedure to obtain the effective coupling, but it does not predict its explicit analytical form in terms of langragian coefficient. In this regard it is useful to use the effective field theory of dark energy to derive general analytical expression for $G^{\Psi}_{\rm{eff}}$, with the advantage of allowing to derive precise predictions for the interconnection with other observables Romano:2025apm; Romano:2025pcs.
The Bardeen potentials are related by
$\Phi-\Psi=\frac{a^{2}}{M^{2}_{\rm{eff}}}\Pi_{DE}\,$, (8)
where $\Pi_{DE}$ is the effective scalar anisotropy potential of dark energy, from which we get
$\nabla^{2}(\Psi+\Phi)=\nabla^{2}\left(2\Psi+\frac{a^{2}}{M^{2}_{\rm{eff}}}\Pi_{DE}\right)=a^{2}G^{\Psi}_{\rm{eff}}\delta\rho_{m}+\frac{a^{2}}{M^{2}_{\rm{eff}}}\nabla^{2}\Pi_{DE}=a^{2}G^{\Psi+\Phi}_{\rm eff}\delta\rho_{m}\,$, (9)
where we have defined
$G^{\Psi+\Phi}_{\rm eff}=G^{\Psi}_{\rm{eff}}+\frac{1}{M^{2}_{\rm{eff}}}\frac{\nabla^{2}\Pi_{DE}}{\delta\hat{\rho}_{m}}=\frac{1}{M^{2}_{\rm{eff}}}\left(1+\frac{\delta\hat{\rho}_{DE}}{\delta\hat{\rho}_{m}}+\frac{\nabla^{2}\Pi_{DE}}{\delta\hat{\rho}_{m}}\right)\,.$ (10)
In momentum space we obtain
$\frac{k^{2}}{a^{2}}(\Psi_{k}+\Phi_{k})=-a^{2}G^{\Psi+\Phi}_{\rm eff}(\eta,k)\delta\rho_{m,k}\,\,$, (11)
where
$G^{\Psi+\Phi}_{\rm eff}(\eta,k)=G^{\Psi}_{\rm{eff}}(\eta,k)-\frac{1}{M^{2}_{\rm{eff}}}\frac{k^{2}\Pi_{DE,k}}{\delta\hat{\rho}_{m,k}}\,.$ (12)
Examples of explicit analytical calculations of $G_{\rm{eff}}^{\Psi+\Phi}$ and $G_{\rm{eff}}^{\Psi}$ for Horndeski theories in the deep sub-horizon limit, can be found in Linder:2015rcz. Note that this definition of effective gravitational coupling is consistent with the scale and time dependent ansatz used in the phenomenological analysis of large scale structure data Ishak:2024jhs.
@@ IV Cosmological effects on gravitational waves
where $h_{+}$ and $h_{\times}$ are the tensor modes, $h_{V1,2}$ the vector polarizations, and $h_{S}$ and $h_{L}$ the transverse and longitudinal scalar modes. Ignoring the interaction of GWs with matter, the free action in a FRW background is given by Romano:2024apw; Gubitosi:2012hu
@@
where $\mathcal{H}=a^{\prime}/a$. Including the interaction terms we get
@@
Note that since cosmological perturbations are space and time dependent, in general also $G^{ij}_{\rm{eff}}(\eta,x)$ is expected to be also space and time dependent. Contrary to GR, in a general modified gravity theory the vector and scalar degrees of freedom cannot be gauged away, because even in absence of matter perturbations, the DE effective energy tensor will act as source for these Flanagan:2005yc. Note that the GWs propagation equation can be always re-cast in the form in eq.(16), by absorbing in $\delta T^{DE}_{ij}$ any modified gravity effect, including for example graviton mass terms deRham:2014zqa. An alternative effective approach consists in encoding the effects of the source term in a frequency-time-polarization dependent effective speed Romano:2022jeh, which can be convenient for the calculation of the effects on gravity modification on the GW-EMW distance ratio.
@@ V General relativity limit
VI Local negative effective gravitational coupling
From eq.(5), we can see that the effective gravitational coupling can be negative when
$\frac{\delta\rho_{DE}}{\delta\rho_{m}}<-1\,.$ (23)
Assuming a positive matter density perturbation $\delta\rho_{m}$, the effective gravitational coupling can be locally negative if the dark energy density perturbations $\delta\rho_{DE}$ are negative. This can happen locally in regions of space where $\delta\rho_{DE}(\eta,x^{i})<-\delta\rho_{m}(\eta,x^{i})$, or in terms of the density contrast, when $\delta_{DE}(\eta,x^{i})<-\delta_{m}(\eta,x^{i})$, i.e. a sufficiently large local dark energy under-density produces a negative effective gravitational coupling. This is a local effect which depends on the local fluctuations of dark energy perturbations, and does not require negative energies, since the total dark energy density can be positive even if $\delta_{DE}<0$. Since the minimum value of the density contrast is by definition 1-1, the effective gravitational coupling can be negative only for $\delta_{m}<1$.
VII Phantom dark energy induced structure suppression
Beside the local effect mentioned above, there can also be a cosmological effect when matter and dark energy perturbations are adiabatic, i.e. when $\delta\rho_{DE}/\rho^{\prime}_{DE}=\delta\rho_{m}/\rho^{\prime}_{m}$, in which case eq.(5) gives
$G^{\Psi}_{\rm{eff}}=\frac{1}{M^{2}_{\rm eff}}\left(1+\frac{\rho^{\prime}_{DE}}{\rho^{\prime}_{m}}\right)=\frac{1}{M^{2}_{\rm eff}}\left[1+(1+w_{DE})\frac{\rho_{DE}}{\rho_{m}}\right]\,$, (24)
where we have used the continuity equation for dark energy, $\rho_{DE}^{\prime}=-3/a\Big[1+w_{DE}(a)\Big]\rho_{DE}$, and for matter, $\rho_{m}^{\prime}=-3/a\,\rho_{m}$. At late time, when dark energy is dominating, and if $w_{DE}<-1$, the effective gravitational coupling can be reduced with respect to its early time value $1/M^{2}_{\rm eff}$, corresponding to when dark energy is subdominat, i.e $\rho_{DE}/\rho_{m}\approx 0$. Quite remarkably, the recent analysis of the Dark Energy Spectroscopic Instrument (DESI) DESI:2024mwx shows a preference for dynamical dark energy with w<1w<-1, and at low redshift there is also an evidence of suppression of structure Chen:2024vuf. In fig.(1) it is shown the redshift dependence of the effective gravitational coupling corresponding to the dark energy equation of state parameters estimated by DESI DESI:2024mwx, confirming that at low redshift there can be a reduction compared to the early time value, and providing an explanation for the observed low redshift structure suppression Chen:2024vuf. This shows that the observed structure suppression at low redshift could be another manifestation of the background dark energy equation of state, providing further independent evidence for this type of dynamical dark energy.
Figure 1: The effective gravitational coupling corresponding to the dark energy equation of state parameters estimated by DESI is plotted as function of redshift, in units of $1/M^{2}_{\rm eff}$. The plot corresponds to $w(a)=w_{0}+w_{a}(1-a)$, with $w_{0}=-0.827$ and $w_{a}=-0.75$, which are the best fit values of the DESI+CMB +PantheonPlus analysis DESI:2024mwx.
VIII Conclusions
For scalar perturbations the scale dependency of the effective gravitational coupling allows to provide an effective description for the screening mechanism, i.e. the suppression of the effects of dark energy on small scales, which is often necessary in order to pass local gravitational tests. The scalar perturbations effective gravitational coupling can be locally negative, in regions with local dark energy under-densities. For adiabatic perturbations the effective gravitational coupling for scalar perturbations can be negative for phantom dark energy, providing a possible explanation for the observed structure suppression at low redshift Chen:2024vuf.
While in this paper we have not considered specific dark energy models, the generality of the assumptions have allowed to derive a general model independent theoretical framework for the study of the effects of dark energy perturbations on cosmological perturbations which can be applied to a large class of theories, and can be convenient for model independent analysis of observational data.
Appendix A Physical implications for scalar and tensor perturbations
We have shown that the effects of dark energy on cosmological perturbations can be modeled by a momentum, time and polarization dependent effective gravitational coupling. A direct implication of this general result is that the gravitational energy radiated by a binary system can acquire an additional dependence on the GWs frequency and polarization. This is due to the fact that the calculation of the radiated power involves the solution of the wave equation for a binary coalescence, and the momentum and polarization dependence of the effective gravitational coupling will induce new effects. Formally we can write the solution of the GWs equation as
$h_{ij}=\Box^{-1}\left({G^{ij}_{\rm{eff}}}\delta T^{m}_{ij}\right)\,$, (25)
which differs from the GR case in which $G^{ij}_{\rm{eff}}$ is a constant. Since $G^{ij}_{\rm{eff}}$ is polarization and space dependent, different polarizations can have a different radiated power. Writing the wave equation in momentum space, the momentum dependence of $G^{ij}_{\rm{eff}}$ would also induce a new additional frequency dependency. In a general modified gravity theory there can be GWs of scalar and vector polarization Ezquiaga:2018btd, which cannot be gauged away, contrary to GR. At leading order we have
$h_{ij}=\frac{2G^{ij}_{\rm{eff}}}{r}\ddot{Q}_{ij},$ (26)
For scalar perturbations the scale dependence of the effective gravitational coupling allows to model phenomena such as the screening mechanism Vainshtein:1972sx, i.e. a suppression the effects of dark energy on small scales, in order to pass local gravitational tests. This allows for example to model modified gravity effects in such a way that they are important only on sufficiently large scales.
@@ Appendix A Physical implications for scalar and tensor perturbations
Appendix B Encoding higher order perturbations effects in the effective coupling
@@
We have so far considered linear order metric perturbations, and at higher order $p$ other gauge invariant variables ${h}^{(p)}_{\mu\nu},\Psi^{(p)},\Phi^{(p)}$ can be defined Mollerach:2003nq; DeLuca:2019ufz; Chang:2020iji. Denoting as $\hat{\Psi}^{(p)}$ and $\delta\hat{\rho}^{(p)}_{m}$ the solutions of the perturbations equations of order p
$E^{(p)}[\delta g^{(p)}_{\mu\nu},\delta\rm T^{(p)}_{\mu\nu}]=0\,$, (28)
we can define the effective dark energy perturbations $\delta\rho^{(p),\rm{eff}}_{DE}$ by
$\nabla^{2}\hat{\Psi}^{(p)}=a^{2}\frac{1}{M^{2}_{\rm{eff}}}\left[\delta{\hat{\rho}}^{(p)}_{m}+\delta\rho^{(p),\rm{eff}}_{DE}\right]\,$, (29)
which shows that the solution $\Psi^{(p)}$ of order p can be obtained as solution of the effective eq.(29) for an appropriate choice of $\delta\rho^{(p),\rm{eff}}_{DE}$. Following a method similar to the one used to derive eq.(4) we can then obtain the effective Poisson’s equation of order p
$\nabla^{2}\Psi^{(p)}=a^{2}G^{\Psi(p)}_{\rm{eff}}\delta{\rho}^{(p)}_{m}\,$,\\ (30)
@@
$G^{\Psi(N)}_{\rm{eff}}=\frac{1}{\delta\rho^{(N)}_{m}}\sum^{N}_{p=1}G^{\Psi(p)}_{\rm{eff}}\delta\rho^{(p)}_{m}\,$. (34)
@@
For GWs we can follow a similar effective approach. First it is convenient to note that eq.(15) can be written in terms of the D’Alambert operator in curved space as
$\Box h^{(1)}=0\,$, (35)
defined by the effective metric Romano:2023bzn; Romano:2022jeh,
$ds_{eff}^{2}=M^{2}_{\rm{eff}}a^{2}\Big[c_{ij}d\eta^{2}-\frac{\delta_{ij}}{c_{ij}}dx^{i}dx^{j}\Big]\,$, (36)
where we are denoting explicitly the order of perturbations. Using a similar procedure to the one outlined above for $\delta\rho^{(p),\rm{eff}}_{DE}$, we can define $\delta T^{DE(p)}_{ij,\rm{eff}}$ by
$\Box\hat{h}^{(p)}_{ij}=\frac{a^{2}}{M^{2}_{\rm{eff}}}\left[\delta\hat{T}^{m(p)}_{ij}+\delta T^{DE(p)}_{ij,\rm{eff}}\right$]\,, (37)
where $\hat{h}$ and $\delta\hat{T}^{m(p)}_{ij}$ are the solutions of the cosmological perturbations equations of order p. This shows that any solution $\hat{h}^{(p)}$ can be obtained as a solution of the effective eq.(37), by appropriately defining $\delta T^{DE(p)}_{ij,\rm{eff}}$. From Eq.(37) we can then obtain the effective equation in terms of the effective gravitational coupling
$\Box h^{(p)}_{ij}=a^{2}{G^{ij(p)}_{\rm{eff}}}\delta T^{m(p)}_{ij}\,$, (38)
[…]
The d’Alambert operator is defined w.r.t. to the same effective metric in all these equations, so they can be summed to give
$\Box h^{(N)}_{ij}=a^{2}{G^{ij(N)}_{\rm{eff}}}\delta T^{m(N)}_{ij}\,$, (40)
where we have used the is linearity of the d’Alambert operator, and defined the summed GWs and energy-tensor perturbations
$h^{(N)}_{ij}=\sum^{N}_{p=1}h^{(p)}_{ij}\quad,\quad T^{(N)m}_{ij}=\sum^{N}_{p=1}\delta T^{(p)m}_{ij}\,$, (41)
and
$G^{ij(N)}_{\rm{eff}}=\frac{1}{\delta T^{(N)m}_{ij}}\sum^{N}_{p=1}{G^{ij(p)}_{\rm{eff}}}\delta T^{(p)m}_{ij}\,.$ (42)